An object of mass 3kg is acted upon by two horizontal forces of 6N and 3N In the opposite directions Calculate the acceleration produced by the system.


1
Expert's answer
2021-02-06T10:04:54-0500

Let's choose the rightwards as the positive direction. Let's assume that the 6 N force directed rightward and the 3 N force directed leftward. Then, we can find the net force acting on the object:


"F_{net}=6\ N-3\ N=3\ N."

Finally, from the Newton's Second Law of Motion we can find acceleration produced by the system:


"a=\dfrac{F}{m}=\dfrac{3\ N}{3\ kg}=1.0\ \dfrac{m}{s^2}."